EXPLICIT IRREDUCIBLE REPRESENTATIONS OF THE IWAHORI-HECKE ALGEBRA OF TYPE F_{4}

ARUN RAM D. E. TAYLOR

Abstract. A general method for computing irreducible representations of Weyl groups and Iwahori-Hecke algebras was introduced earlier by the first author. The representations of the algebras of types A_{n}, B_{n}, D_{n} and G_{2} were computed and it is the purpose of this paper to extend these computations to F_{4}. The main goal here is to compute irreducible representations of the Iwahori-Hecke algebra of type F_{4} by only using information in the character table of the Weyl group.